1. General status updates/Schedule review.
Consider day change to Wednesday to include Rick.
Eric -- a lot of work with Gianni to simulate → trigger filters → logger … found bugs in spawning more than one trigger process in an Event Builder. Event numbering was confused, now uses Event Window Tag .. but in simulation mode event window tags repeat.. Still some issues. Yesterday daq12 (which is the DCS dev node) locked up -- similar to previous DCS lockup.
Ron --
Greg --
Monica -- Found stolen monitor. DTC in daq09 is dead, worked with Rick to understand. Daq05 working differently than daq09 - daq09 is worse. Will come next Tuesday to work with Rick again.
Antonio -- Working on export/import with Pheobus and ots DCS web GUI. Works, but working on details - should be interchangeable.
Craig --
Gary --
Bertrand -- Working on improving timing performance of some trigger modules. Giani showed ~6ms, and some opportunities for improvement - maybe need art team.
Glenn -- Vivian’s meeting is good. Working on Mu2e CRV test stand. Will start looking at DCS soon. Dcs01 and dcs02 are managed by SLAM. Need newer Java.
Lei -- Sten will continue working with Rick next week.
Vivian -- Giani made snapshot of his system. Anyone with access mu2edaq should be able to clone the snapshot.

2. Status of DTC to ROC interface.
3. Status of EPICS on DAQ cluster.
4. Status of core ROC firmware testing.